Evaluate the following integrals:
(i) ∫ x ln(x + 1)dx

(ii) ∫ ((sin³ (In x) cos²(In x)) / (x)) dx
1
Expert's answer
2020-07-01T15:30:54-0400

i. Let "I=\int xln[1+x]dx"

Using integration by parts : "I=ln[1+x]\int xdx - \int[(d[ln[1+x]/dx) * \int xdx]dx"

"I=ln[1+x]*(x^2/2) -\int[(1/(1+x)*1*(x^2/2)]dx"

or, "I=(x^2/2)*ln[1+x] - (1/2)*\int" "[x^2\/(1+x)]dx"

Let us calculate "\int [x^2/(1+x)]dx"

"=\int(x^2-1+1)/(1+x)dx"

"=\int[(x-1)(x+1)/(x+1)+(1/(1+x))]dx"

"=\int[(x-1) +(1/(1+x))]dx"

"=\int(x-1)dx+\int1/" "(1+x)dx"

"Let 1+x" ="z"

therefore, "dx=dz"

therefore , "I=\int(x-1)dx+\int(1/z)dz"

"=(x^2\/2)-x+lnz"

Putting "z=1+x"

"I=(x^2\/2)-x+ln[1+x]"

Putting this back to "I" we get:

"I=(x^2\/2)ln[1+x] -(1\/2)*[x^2\/2-x+ln[1+x]]+C"

"=(x^2\/2)ln[1+x]-(x^2\/4)+(x\/2)-(1\/2)*ln[1+x]+C"

where C is an integration constant (Answers)




ii. Let "I=\int[sin^3(lnx)cos^2(lnx)]/xdx"

Let "lnx=z"

therefore, "(1\/x)dx=dz"

therefore, "I=\int[sin^3zcos^2z]dz"

"=\int[sin^2zcos^2sinz]dz"

Let "cosz=u"

"-sinzdz=du"

therefore, "I=\int-(1-u^2)*u^2du"

"=\int(-u^2+u^4)du"

"=-u^3\/3+u^5\/5 +c" where C is an integration constant

Putting back "u=cosz" we get

"I=-cos^3z\/3+cos^5z\/5 + C"

Again putting back "z=lnx" we get:

"I=-cos^3(lnx)\/3 + cos^5(lnx)\/5 +C (Answer)"